# Polynomial bounds for surfaces in cusped 3-manifolds

### v2

@@ 5. Placing curves on the chunk boundary
Any two consecutive intersections in a combination will be connected by an arc of $\partial Z_{i}$ in a face of a chunk. Because each face of a chunk is a disk, such an arc is unique up to isotopy. The only way the boundary component might not be unique is if multiple arcs with intersections on the same edges interleave in a face in different ways. But for a fixed combination, there will be a unique way to connect such arcs into a closed curve.
Therefore, a combination determines a closed curve on the chunk boundary. If we have two identical combinations, they will determine parallel curves on the chunk boundary. These are isotopic on the chunk boundary, as required. ∎
@@
Counting combinations gives us the following lemma, which is similar to part of the argument of Hass, Thompson, and Tsvietkova [15, 16], despite the fact that here the 3-manifold is more general, and the normal subsurfaces $Z_{i}$ are general surfaces with boundary, while in [15, 16] analogous surfaces are disks. Also, unlike in [16], a surface $Z$ with non-meridianal boundary is not necessarily a spanning surface for a link, i.e. its boundary curve does not have to have intersection number 11 11 with a meridian.
@@
By Lemma 4.3, there are at most $-20\chi(Z)$ intersections with chunk edges for $\gamma$, as long as $\gamma$ is not the boundary of a $BBBB$ or $BBSS$ disk. But even if it is the boundary of a $BBBB$ or $BBSS$ disk, there are only $4<-20\chi(Z)$ intersections. Thus there are at most $(6n)^{-20\chi(Z)}$ combinations of length $-20\chi(Z)$ on $\Pi^{+}$. However, we wish to count curves with at most $-20\chi(Z)$ intersections, meaning we need to count ones with $-20\chi(Z)$ intersections, with $-20\chi(Z)-1$ intersections, etc., down to length 22 22. For this, we make use of the observation that our count of the combinations is already an overcount: For a combination with $s<-20\chi(Z)$ intersections, there will be a ‘last’ intersection (depending on our choice of starting intersection). Define a combination of length $-20\chi(Z)$ associated to the shorter combination by requiring the combination to hit the same last edge consecutively over and over. Such a combination does not correspond to a curve of $\partial Z_{i}$, but is counted in the bound $(6n)^{-20\chi(Z)}$. Therefore, $(6n)^{-20\chi(Z)}$ is an upper bound for the number of curves on $\Pi^{+}$. ∎
@@ 6. Surfaces in classical alternating link complements
In prior sections, we worked with a link $L$ projected onto a surface $\Pi$ in a general 3-manifold $Y$ satisfying a few mild hypotheses, and developed machinery for weakly generalised alternating links in $Y$. In this section, we temporarily restrict our attention to $Y=S^{3}$ and classical alternating links on $S^{2}$.
@@
Suppose $Z_{i}$, $i=1,\dots,k$, are all normal subsurfaces of an essential surface $Z$ in chunks that are not $BBBB$ or $BBSS$ disks. By Lemma 5.4, at most $C(n,\chi(Z))=(6n)^{-20\chi(Z)}$ curves on the boundary of the chunk could be a boundary component of $Z_{i}$. Because each $Z_{i}$ is a disk, $C(n,\chi(Z))$ therefore gives a bound on the number of possibilities for the disks $Z_{i}$. Set aside all $Z_{i}$ that are $BBBB$ and $BBSS$ disks. The number of options for all remaining $\partial Z_{i}$ to be placed on the boundaries of the chunks is at most $C(n,\chi(Z))$. By Theorem 2.8, these surfaces uniquely determine the surface $Z$. By Lemma 4.1, there are at most $-4\chi(Z)$ such surfaces.
Therefore, $C(n,\chi(Z))^{-4\chi(Z)}=(6n)^{80\chi(Z)^{2}}$ is an upper bound for the number of such surfaces. ∎
@@
To obtain the original $Z$, we need to tube back together all boundary components created by meridianal compressions. Hence we need to multiply the upper bound by the number of ways to do these tubings. Mossessian studied ways to tube together surfaces in [33]. Although that paper is concerned with closed Heegaard surfaces, the argument in [33, Lemma 3.7] applies in more generality to construct tubed surfaces. It shows that a tubed surface is determined by any $-2\chi(Z)+1$ element subset of the $-4\chi(Z)+2$ boundary components of $Z^{\prime}$ to be tubed. Thus there are at most (${-4\chi(Z)+2\choose-2\chi(Z)+1}$ such tubings. The final bound follows from the fact that (${n\choose k}\leq 2^{n}$ for $k\leq n$. ∎
@@
With the same hypotheses on $L$ and $\pi(L)$, the bound in Theorem 6.1 holds if one considers the number of isotopy classes of all properly embedded connected essential surfaces of fixed Euler characteristic $\chi$, instead of just one such fixed topological surface.
@@ 7. The number of surfaces in arbitrary 3-manifolds
In Theorem 6.1, each subsurface $Z_{i}$ is a disk in a chunk that is a ball, and therefore $Z_{i}$ is uniquely determined by its boundary. More generally, we will have subsurfaces $Z_{i}$ that can have varying genera. To count the total number of embedded non-isotopic surfaces in such cases, we need a bound on the number of incompressible surfaces embedded in the chunk $C$ with fixed boundary curves and with a fixed genus, both orientable and nonorientable. This will be denoted by $X$. Assuming we have such a bound, we obtain a bound on the number surfaces embedded in a weakly generalized alternating link exterior, as follows.
@@
Let $L$ be a weakly generalized alternating link in $Y$, with projection $\pi(L)$ with $n$ crossings on $\Pi$, where $Y$, $\Pi$ and $L$ satisfy Assumptions 2.1 and 2.9. Suppose that for each 3-manifold component $\Sigma$ of $Y-N(\Pi)$ there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in $\Sigma$ with fixed genus and fixed boundary curves on $\partial N(\Pi)\cap\partial\Sigma$.
Fix a topological surface $Z$, with genus $g$, Euler characteristic $\chi$, that is either orientable or non-orientable. Then the number of ways $(Z,\partial Z)$ can be properly embedded in $(Y-N(L),\partial N(L))$ as an essential, meridianally incompressible surface, up to isotopy, is at most:
@@
We do not include a result equivalent to Corollaries 6.2 and 6.4 in this section. The proofs of theorems in this section still construct (and therefore count) any surface of given genus, Euler characteristic and orientability type. But due to the classification of surfaces, once the genus, Euler characteristic and orientability type is fixed, the surface is unique topologically. We are only obtaining embeddings of this surface up to isotopy. We could have reformulated the above theorem as giving the number of non-isotopic embedded surfaces of fixed genus, Euler characteristic and orientability type.
@@
Recall that an embedded, meridianally incompressible essential surface $Z$ consists of subsurfaces $Z_{i}$ in chunks. As in Theorem 6.1, each $Z_{i}$ that is not a $BBBB$ or $BBSS$ disk can have from 00 0 to $-10\chi(Z)$ boundary components by Lemma 4.3. Each $Z_{i}$ can also have genus or nonorientable genus from 00 0 to $g$. Moreover, for a fixed collection of boundary components and fixed $g$, $Z_{i}$ may be one of at most $X$ subsurfaces by assumption.
Suppose first the boundary of $Z_{i}$ consists of $w$ curves already placed on the boundary of chunks. We then have at most $X$ options for $Z_{i}$ of genus $g^{\prime}$ inside its chunk. Since the genus of $Z_{i}$ may be from 00 0 to $g$, we have
@@
To obtain the original closed surface $Z$ from $Z^{\prime}$, we need to tube together the punctures introduced by meridian compression. Hence we need to multiply the upper bound for the number of meridianally incompressible surfaces by the number of ways to tube its meridianal punctures. As in the proof of Theorem 6.3, by work of Mossessian [33, Lemma 3.7], the number of tubings producing non-isotopic embedded surfaces is:
(${-4\chi(Z)+2\choose-2\chi(Z)+1}$
Hence the bound is:
$\displaystyle X^{-4\chi}\cdot(g+1)^{-4\chi+1}\cdot(6n)^{-800\chi^{3}+80\chi^{2}}{-4\chi+2\choose-2\chi+1}$
@@
$\displaystyle\leq(2X(g+1))^{-4\chi+2}\cdot(6n)^{-800\chi^{3}+80\chi^{2}}$
since (${n\choose k}\leq 2^{n}$. ∎
@@ 8.1. Surfaces in thickened tori and solid tori
The following is stated by Przytycki [34, Theorem 2.3 and Corollary 2.5].
@@
Let $F$ be a properly embedded incompressible surface in $T^{2}\times I$ that is not a boundary parallel disk. Then either
@@
    $F$ is isotopic to an annulus of type $(\gamma)\times I$ for a nontrivial simple closed curve $\gamma$ in $T^{2}$, or
@@
    $F$ is an annulus or torus parallel to the boundary, or
@@
    $F$ is isotopic to a nonorientable surface uniquely determined by two different slopes, one on $T^{2}\times\{0\}$ and one on $T^{2}\times\{1\}$.
@@
We may isotope the surface $F$ so that it meets the core of the solid torus transversely in a finite number of points; equivalently, $F$ meets a regular neighbhourhood $V$ of the core in a finite number of meridian disks. Then $G=F\cap(S^{1}\times D^{2}-V)$ is an incompressible surface, possibly with multiple components, in a thickened torus. By Theorem 8.1, each component of $G$ has one of three forms unless it is a boundary parallel disk.
@@
The following result is more concrete than Theorems 7.4 and 7.1, since it does not depend on $X$. Remark 7.3 still applies here, i.e. the following corollary gives the bound on the number of embedded non-isotopic surfaces with fixed genus (orientable or non-orientable), Euler characteristic and orientability type. Topologically they are all embeddings of the same surface.
@@
Suppose that $\Pi$ is a torus in the 3-manifold $Y$ having one of the following forms:
@@
Let $L$ be a weakly generalized alternating link in $Y$ with projection $\pi(L)$ on $\Pi$, with a cellular diagram and in cases (a) and (c), representativity $r(\pi(L),\Pi)>4$. Let $n$ be the number of crossings in the diagram $\pi(L)$. Let $Z$ be a fixed connected topological surface with Euler characteristic $\chi$. Then up to isotopy, the number of ways that $(Z,\partial Z)$ can be properly embedded as an essential, meridianally incompressible surface in $(Y-N(L),\partial N(L))$ is at most:
@@
Observe that $Y$, $\Pi$, and $L$ above satisfy Assumptions 2.1 and 2.9, so Theorem 7.1 and Theorem 7.4 apply.
In all cases, the chunks consist of solid tori or thickened tori. When we isotope a properly embedded essential surface homeomorphic to $Z$ into normal form, it will have boundary only on the torus corresponding to $\partial N(\Pi)$. Hence Theorem 8.1 (in cases (b) and (c)) and Corollary 8.3 (in case (a) and (c)) apply. Therefore genus and boundary curves on $\partial N(\Pi)$ uniquely determine the surface up to isotopy if $Z$ is nonorientable. Theorem 8.1 also allows $Z$ to be a boundary parallel annulus, i.e. $Z$ is then parallel to the torus $\Pi$. There are at most two different boundary parallel annuli with the same boundary components. Thus for all these manifolds $X\leq 2$, and we may set $X=2$ in Theorem 7.1 and Theorem 7.4. ∎

### v3

@@ 5. Placing curves on the chunk boundary
Any two consecutive intersections in a combination will be connected by an arc of $\partial Z_{i}$ in a face of a chunk. Because each face of a chunk is a disk, such an arc is unique up to isotopy. The only way the boundary component might not be unique is if multiple arcs with intersections on the same edges interleave in a face in different ways. But for a fixed combination, there will be a unique way to connect such arcs into a closed curve. Therefore, a combination determines a closed curve on the chunk boundary. If we have two identical combinations, they will determine parallel curves on the chunk boundary. These are isotopic on the chunk boundary, as required. ∎
@@
Counting combinations gives us the following lemma, which is similar to part of the argument of Hass, Thompson, and Tsvietkova [24, 25], despite the fact that here the 3-manifold is more general, and the normal subsurfaces $Z_{i}$ are general surfaces with boundary, while in [24, 25] analogous surfaces are disks. Also, unlike in [25], a surface $Z$ with non-meridianal boundary is not necessarily a spanning surface for a link, i.e. its boundary curve does not have to have intersection number $1$ with a meridian.
@@
By Lemma 4.3, there are at most $-20\chi(Z)$ intersections with chunk edges for $\gamma$, as long as $\gamma$ is not the boundary of a $BBBB$ or $BBSS$ disk. But even if it is the boundary of a $BBBB$ or $BBSS$ disk, there are only $4<-20\chi(Z)$ intersections. Thus there are at most $(6n)^{-20\chi(Z)}$ combinations of length $-20\chi(Z)$ on $\Pi^{+}$. However, we wish to count curves with at most $-20\chi(Z)$ intersections, meaning we need to count ones with $-20\chi(Z)$ intersections, with $-20\chi(Z)-1$ intersections, etc., down to length $2$. For this, we make use of the observation that our count of the combinations is already an overcount: For a combination with $s<-20\chi(Z)$ intersections, there will be a ‘last’ intersection (depending on our choice of starting intersection). Define a combination of length $-20\chi(Z)$ associated to the shorter combination by requiring the combination to hit the same last edge consecutively over and over. Such a combination does not correspond to a curve of $\partial Z_{i}$, but is counted in the bound $(6n)^{-20\chi(Z)}$. Therefore, $(6n)^{-20\chi(Z)}$ is an upper bound for the number of curves on $\Pi^{+}$. ∎
@@ 6. Surfaces in classical alternating link complements
In prior sections, we worked with a link $L$ projected onto a surface $\Pi$ in a general 3-manifold $Y$ satisfying a few mild hypotheses, and developed machinery for weakly generalized alternating links in $Y$. In this section, we restrict our attention to $Y=S^{3}$ and classical alternating links on $S^{2}$.
@@
Suppose $Z_{i}$, $i=1,\dots,k$, are all normal subsurfaces of an essential surface $Z$ in chunks that are not $BBBB$ or $BBSS$ disks. By Lemma 5.4, at most $C(n,\chi(Z))=(6n)^{-20\chi(Z)}$ curves on the boundary of the chunk could be a boundary component of $Z_{i}$. Because each $Z_{i}$ is a disk, $C(n,\chi(Z))$ therefore gives a bound on the number of possibilities for the disks $Z_{i}$. Set aside all $Z_{i}$ that are $BBBB$ and $BBSS$ disks. The number of options for all remaining $\partial Z_{i}$ to be placed on the boundaries of the chunks is at most $C(n,\chi(Z))$. By Theorem 2.8, these surfaces uniquely determine the surface $Z$. By Lemma 4.1, there are at most $-4\chi(Z)$ such surfaces. Therefore, $C(n,\chi(Z))^{-4\chi(Z)}=(6n)^{80\chi(Z)^{2}}$ is an upper bound for the number of such surfaces. ∎
@@
Corollary 6.3.
Let $\pi(L)$ be a prime alternating projection of a link $L$ onto $S^{2}$ in $Y=S^{3}$, with $n$ crossings, and suppose that $S^{3}-L$ is hyperbolic. Then for any fixed connected topological surface $Z$ with Euler characteristic $\chi$, the number of ways, up to isotopy, that $Z$ can be properly embedded in $S^{3}-N(L)$ as a quasifuchsian surface is at most $(6n)^{80\chi(Z)^{2}}$.
Proof.
Essential surfaces in hyperbolic 3-manifolds are exactly one of the following three types (see [8]): quasifuchsian, a virtual fiber, contain accidental parabolics. Since a meridianal compression gives an accidental parabolic, any quasifuchsian surface satisfies the hypothesis of Theorem 6.1. ∎
@@
To obtain the original $Z$, we need to tube back together all boundary components created by meridianal compressions. Hence we need to multiply the upper bound by the number of ways to do these tubings. Mossessian studied ways to tube together surfaces in [46]. Although that paper is concerned with closed Heegaard surfaces, the argument in [46, Lemma 3.7] applies in more generality to construct tubed surfaces. It shows that a tubed surface is determined by any $-2\chi(Z)+1$ element subset of the $-4\chi(Z)+2$ boundary components of $Z^{\prime}$ to be tubed. Thus there are at most ${-4\chi(Z)+2\choose-2\chi(Z)+1}$ such tubings. The final bound follows from the fact that ${n\choose k}\leq 2^{n}$ for $k\leq n$. ∎
@@
With the same hypotheses on $L$ and $\pi(L)$, the bound in Theorem 6.1 holds if one considers the number of isotopy classes of all properly embedded connected essential surfaces of fixed Euler characteristic $\chi$, instead of just one such fixed surface.
@@ 7. The number of surfaces in arbitrary 3-manifolds
In Theorem 6.1, each subsurface $Z_{i}$ is a disk in a chunk that is a ball, and therefore $Z_{i}$ is uniquely determined by its boundary. More generally, we will have subsurfaces $Z_{i}$ that can have varying genera. To count the total number of embedded non-isotopic surfaces in such cases, we need a bound on the number of incompressible surfaces embedded in the chunk $C$ with fixed boundary curves and with a fixed genus, both orientable and nonorientable. This will be denoted by $X$. Assuming we have such a bound, we obtain a bound on the number surfaces embedded in a weakly generalized alternating link exterior.
@@
Let $L$ be a weakly generalized alternating link in $Y$, with projection $\pi(L)$ with $n$ crossings on $\Pi$, where $Y$, $\Pi$ and $L$ satisfy Assumptions 2.1 and 2.9. Suppose that for each 3-manifold component $\Sigma$ of $Y-N(\Pi)$ there is a universal bound $X$ on the number of isotopy classes of incompressible surfaces properly embedded in $\Sigma$ with fixed genus and fixed boundary curves on $\partial N(\Pi)\cap\partial\Sigma$. Fix a topological surface $Z$, with genus $g$, Euler characteristic $\chi$, that is either orientable or non-orientable. Then the number of ways $(Z,\partial Z)$ can be properly embedded in $(Y-N(L),\partial N(L))$ as an essential, meridianally incompressible surface, up to isotopy, is at most:
@@
We do not include a result equivalent to Corollaries 6.2 and 6.5 in this section. The proofs of theorems in this section still construct (and therefore count) any surface of given genus, Euler characteristic and orientability type. But due to the classification of surfaces, once the genus, Euler characteristic and orientability type is fixed, the surface is unique topologically. We are only obtaining embeddings of this surface up to isotopy. We could have reformulated the above theorem as giving the number of non-isotopic embedded surfaces of fixed genus, Euler characteristic and orientability type.
@@
Recall that an embedded, meridianally incompressible essential surface $Z$ consists of subsurfaces $Z_{i}$ in chunks. Each $Z_{i}$ that is not a $BBBB$ or $BBSS$ disk may have from $0$ to $-10\chi(Z)$ boundary components by Lemma 4.3. Each $Z_{i}$ may also have genus or nonorientable genus from $0$ to $g$. And for a fixed collection of boundary components and fixed $g$, $Z_{i}$ may be one of at most $X$ subsurfaces by assumption.
Suppose first the boundary of $Z_{i}$ consists of $w$ curves already placed on the boundary of chunks. We then have at most $X$ options for $Z_{i}$ of genus $g^{\prime}$ inside its chunk. Since the genus of $Z_{i}$ may be from $0$ to $g$, we have
@@
Corollary 7.4.
With all the hypotheses of Theorem 7.1, suppose in addition that $Y-L$ admits a hyperbolic structure. Then the bound of Theorem 7.1 bounds the number of ways, up to isotopy, that $Z$ can be properly embedded as a quasifuchsian surface.
Proof.
As before, a quasifuchsian surface is meridianally incompressible. ∎
@@
To obtain the original closed surface $Z$ from $Z^{\prime}$, we need to tube together the punctures introduced by meridian compression. Hence we need to multiply the upper bound for the number of meridianally incompressible surfaces by the number of ways to tube its meridianal punctures. As in the proof of Theorem 6.4, by work of Mossessian [46, Lemma 3.7], the number of tubings producing non-isotopic embedded surfaces is:
$4\chi(Z)+2\choose-2\chi(Z)+1$}
Hence, using the fact that ${n\choose k}\leq 2^{n}$, the bound is:
$\displaystyle X^{-4\chi}$ $\displaystyle\cdot(g+1)^{-4\chi+1}\cdot(6n)^{-800\chi^{3}+80\chi^{2}}{-4\chi+2\choose-2\chi+1}$
@@
$\displaystyle\leq(2X(g+1))^{-4\chi+2}\cdot(6n)^{-800\chi^{3}+80\chi^{2}}\qed$
@@ 8.1. Surfaces in thickened tori and solid tori
The following is stated by Przytycki [47, Theorem 2.3 and Corollary 2.5].
@@
Let $F$ be a properly embedded incompressible surface in $T^{2}\times I$ that is not a boundary parallel disk. Then $F$ is isotopic to either
@@
    an annulus $(\gamma)\times I$, for $\gamma\subset T^{2}$ a nontrivial simple closed curve, or
@@
    an annulus or torus parallel to the boundary, or
@@
    a nonorientable surface that is uniquely determined by two different slopes, one on $T^{2}\times\{0\}$ and one on $T^{2}\times\{1\}$.
@@
We may isotope the surface $F$ so that it meets the core of the solid torus transversely in a finite number of points; equivalently, $F$ meets a regular neighbhourhood $V$ of the core in a finite number of meridian disks. Then $G=F\cap(S^{1}\times D^{2}-V)$ is an incompressible surface, possibly with multiple components, in a thickened torus. By Theorem 8.1, each component of $G$ is a boundary parallel disk or has one of three forms.
@@
The following result is more concrete than Theorems 7.5 and 7.1, since it does not depend on $X$. Remark 7.3 still applies here, i.e. the following corollary gives the bound on the number of embedded non-isotopic surfaces with fixed genus (orientable or non-orientable), Euler characteristic and orientability type.
@@
Suppose that $\Pi$ is a torus in the 3-manifold $Y$ that is either:
@@
Let $L$ be a weakly generalized alternating link in $Y$ with projection $\pi(L)$ on $\Pi$, with a cellular diagram and in cases (a) and (c), representativity $r(\pi(L),\Pi)>4$. Let $n$ be the number of crossings in the diagram $\pi(L)$. Let $Z$ be a fixed connected topological surface with Euler characteristic $\chi$. Then up to isotopy, the number of ways that $(Z,\partial Z)$ can be properly embedded in $(Y-N(L),\partial N(L))$ as an essential, meridianally incompressible surface is at most:
@@
Observe that $Y$, $\Pi$, and $L$ above satisfy Assumptions 2.1 and 2.9, so Theorem 7.1 and Theorem 7.5 apply. In all cases, the chunks consist of solid tori or thickened tori. When we isotope a properly embedded essential surface homeomorphic to $Z$ into normal form, it will have boundary only on the torus corresponding to $\partial N(\Pi)$. Hence Theorem 8.1 (in cases (b) and (c)) and Corollary 8.3 (in case (a) and (c)) apply. Therefore genus and boundary curves on $\partial N(\Pi)$ uniquely determine the surface up to isotopy if $Z$ is nonorientable. Theorem 8.1 also allows $Z$ to be a boundary parallel annulus, i.e. $Z$ is then parallel to the torus $\Pi$. There are at most two different boundary parallel annuli with the same boundary components. Thus for all these manifolds $X\leq 2$, and we may set $X=2$ in Theorem 7.1 and Theorem 7.5. ∎